### v1

# Personalized Federated Learning under Model Dissimilarity Constraints

Abstract
One of the defining challenges in federated learning is that of statistical heterogeneity among clients. We address this problem with Karula, a regularized strategy for personalized federated learning, which constrains the pairwise model dissimilarities between clients based on the difference in their distributions, as measured by a surrogate for the 1-Wasserstein distance adapted for the federated setting. This allows the strategy to adapt to highly complex interrelations between clients, that e.g., clustered approaches fail to capture. We propose an inexact projected stochastic gradient algorithm to solve the constrained problem that the strategy defines, and show theoretically that it converges with smooth, possibly non-convex losses to a neighborhood of a stationary point with rate $O(1/K)$. We demonstrate the effectiveness of Karula on synthetic and real federated data sets.
1 Introduction
Federated learning (FL) is a learning paradigm in which models are trained on data owned by clients in a network which is connected to a central server. As a response to the increasing presence of private and non-centrally stored data, e.g., data stored on mobile devices or hospital databases, the goal of federated learning is to learn models in a collaborative and communication-efficient manner, without exchanging client data. Federated averaging (McMahan et al.,, 2017) is an example of a FL algorithm, in which clients locally perform gradient steps on a global model, and send updated model parameters to the server, which then averages and returns the updated global model to the clients. There are many examples where FL has been applied successfully, leading to greater communication-efficiency and data privacy, e.g., in mobile keyboard prediciton (Hard et al.,, 2018), autonomous driving (Nguyen et al.,, 2022), and healthcare (Qayyum et al.,, 2022). However, FL faces several challenges that arise due to client hetereogeneity, be it in terms of models, hardware, or data. In the case of data heterogeneity, the client data follow different statistical distributions and may have varying number of observations. This means that the performance of standard FL systems, which learn a single global model, can suffer significantly, both in terms of convergence and generalization. At the same time, due to limited data, non-collaboratively training separate models also leads to poor generalization.
Therefore, in settings where data heterogeneity is significant, federated learning systems should ideally learn personalized models in a collaborative manner where each model borrows strength from somewhere else. Personalized federated learning (PFL) strategies attempt to do just that. We will divide these strategies into two categories: (1) strategies that learn local models that borrow strength from a global model; (2) strategies that learn local models that borrow strength from clients identified as similar at the server-level. We refer to the two categories as globally anchored personalization and peer-based personalization, respectively. Globally anchored strategies are appropriate in applications where we expect each client to have limited dissimilarity with any other client in the network, such as in next-word prediction on mobile devices. However, in applications such as recommender systems, distributional differences between two given clients can be very large. Peer-based methods address this issue by attempting to only learn from clients that are similar. One approach to peer-based personalization is to cluster clients, prior to training or concurrently. The cluster model of statistical heterogeneity does however fail to capture more complex overlaps between clients. As an example, in a music recommendation application, suppose one user (client) is only interested in jazz and classical, another in jazz and funk, and yet another in funk and rock. Clearly, no hard clustering of these users will capture these relationships accurately. Other approaches to peer-based personalization allow for more complex relationships between the local models of clients, but to the best of our knowledge no prior work incorporates distributional similarity measures in a principled manner.
Motivated by this, we develop a PFL strategy that uses approximations of the Wasserstein distance to measure distributional similarity between clients, and in turn, leverages similarity by constraining the dissimilarities between the personalized models. Our method approximates the Wasserstein distances in a manner that does not necessitate the exchanging of client data.
1.1 Contributions
The main contributions of this work are three-fold, and are as follows:
  • We propose a peer-based PFL strategy which learns personalized models that are centerally constrained to have limited pairwise dissimilarity. The strategy produces the limits on the model dissimilarities in a principled and completely data-driven manner using approximations of the Wasserstein distances between the empirical client distributions, motivated by forthcoming theory presented in this work.
  • We propose an efficient inexact projected stochastic gradient method for the sceme is presented, which employs variance reduction to deal with partial client participation in each communication round. Due to this variance reduction, we are able to show convergence for smooth, possibly non-convex loss at rate $O(1/K)$ to an $\epsilon$-stationary set of personalized models, where the error $\epsilon$ arises only due to stochastic gradients from the clients and inexact projection at the server level.
  • We conduct numerical studies on synthetic and real heterogeneous data sets with regression and classification tasks, and compare our strategy with a number of other FL and PFL strategies.
1.2 Outline
In §2 we describe some of the related work, relating our contribution to existing literature. Then, in §3, we introduce or proposed strategy for personalized federated learning, which is followed by a convergence analysis in §4. We present numerical experiments validating our proposed method in §5. We conclude by summarizing and discussing our findings in §6.
2 Related work
Statistical heterogeneity presents significant challenges to both traditional and distributed learning systems, affecting inference and generalization. It arises in various applications where data can be assumed to be drawn from distinct distributions. For example, in meta-analysis, studies with participants from different populations are aggregated (Melsen et al.,, 2014); in medicine, different clinical features in patients affect individual health outcomes (Machado et al.,, 2011); and in finance, markets adopt different regimes over time (Ang and Timmermann,, 2012). In the context of FL, the statistical heterogeneity among clients is generally considered one of the defining aspects of the field. The works of Zhao et al., (2018) and Karimireddy et al., (2020) show that both convergence and generalization of conventional FL algorithms are severely affected by statistical heterogeneity. The heterogeneity introduces difficulties in optimizing the global model, as updates from different clients may conflict, reducing the effectiveness of aggregation methods and slowing down convergence. The former work additionally shows that the 1-Wasserstein distance between client distributions can be used to quantify this drift.
One approach to address the challenges caused by statistical heterogeneity is stratification, whereby different models are fit for every value of some categorical parameter, which are assumed to correspond to the different underlying distributions. Stratified models allow each subgroup to be modeled independently, capturing the unique characteristics of each distribution and reducing the bias that arise from pooling heterogeneous data. In clinical trials, it is common to do this by grouping participants into different strata based on relevant clinical features, and fitting a model for each strata (Kernan et al.,, 1999). Stratification does however reduce the number of observations for each model, meaning viability of naive stratification is limited in many settings. PFL models, which adapt to individual clients rather than enforce a single global model, are a special case of stratified models where the stratification is performed at the level of individual clients.
Globally anchored personalization.
A common approach to PFL is to learn local models that are, in some sense, close to a shared global model. Fallah et al., (2020) take this approach using meta-learning to formulate a method that learns a global model, from which each client takes a few local gradient descent steps to produce the local model. Additionally, they show a connection between the convergence of their method and the statistical heterogeneity measured in total variation and Wasserstein distance, respectively. Taking a regularization approach to the problem, the works of Dinh et al., (2020) and Li et al., (2021) propose methods that penalize the squared norm difference between the local and the global model. Yet another approach is given by Deng et al., (2020), whose algorithm learns personalized models that are convex combinations of a shared global model and a completely local model.
Peer-based personalization.
A more flexible personalization strategy is to identify clients that are similar and borrow strength from them directly rather than from a single aggregated model. One way to do this is to cluster the clients. The PFL method proposed by Ghosh et al., (2020) does this iteratively by alternating between optimizing the cluster models and clustering the clients, which is done by finding the cluster whose model achieves the smallest local loss. Regularized methods for peer-based PFL include convex clustering via a group-lasso penalty on the differences of models (Armacki et al.,, 2022) and Laplacian regularization (Dinh et al.,, 2022; Hashemi et al.,, 2024). The empirical results of Dinh et al., show that in fact Laplacian regularization with randomized weights outperform many globally anchored personalization strategies in several tasks.
Similarity measures.
Measures of task and distribution similarity are of great interest in several learning contexts, particularly transfer learning (Zhong et al.,, 2018; Pan,, 2020), multi-task learning (Shui et al.,, 2019), and meta-learning (Zhou et al.,, 2021). These measures quantify how closely related different tasks or data distributions are, influencing decisions on model adaptation, aggregation, and generalization. Common approaches include statistical distance metrics such as the Kullback-Leibler divergence (Kullback and Leibler,, 1951), Jensen-Shannon divergence (Lin,, 1991), and Wasserstein distance (Kantorovich,, 1960). In addition, similarity can be assessed through feature-space representations using methods like cosine similarity or kernel-based measures.
3 Proposed method
In this section, we introduce Karula [^1] The name is a portmanteau of Kantorovich, Rubenstein and Laplace., a strategy for federated learning of personalized models where model differences are constrained by the distributional similarity of client data. We also propose an efficient projected stochastic gradient method with variance reduction to handle partial client participation during communication rounds. This method is also suitable for use in other FL strategies.
3.1 Main idea
Consider the FL setting with $N$ clients in a network that can communicate with a central server. To each client $i\in\{1,\dots,n\}$, we associate an underlying distribution $\mu_{i}\colon 2^{\mathcal{X}}\to{\mbox{\bf R}}_{+}$, with $\mathcal{X}$ being the feature space, and a data set $\mathcal{D}_{i}$ of $N_{i}$ independent samples drawn from this distribution. Ideally, we would learn models $\theta_{i}\in\Theta$ that minimize $f_{i}(\theta_{i})={\mathbbmss{E}}_{\xi\sim\mu_{i}}[\ell(\theta_{i},\xi)]$, where $\ell\colon\Theta\times\mathcal{X}\to{\mbox{\bf R}}$ is a loss function, and $\Theta$ is the parameter space. However, due to limited available data on each client, attempting to directly minimize $f_{i}(\theta_{i})$ leads to poor generalization, so our strategy regularizes the models by approximately solving the constrained problem
$\begin{array}[]{ll}\underset{\theta_{i}\in\Theta}{\mbox{minimize}}&\frac{1}{n}\sum_{i=1}^{n}\alpha_{i}f_{i}(\theta_{i})\\ \mbox{subject to}&\|\theta_{i}-\theta_{j}\|_{2}^{2}\leq tD_{ij},\quad i,j\in\{1,\dots,n\}.\end{array}$ (1)
Here $t\geq 0$ is a regularization parameter, each $\alpha_{i}$ is a positive weight, and each $D_{ij}\geq 0$ is a dissimilarity parameter which quantifies the dissimilarity between the $i$th and $j$th client, and are calculated before solving the minimization problem. This formulation generalizes the standard FL formulation, since by selecting $t=0$ the problem is equivalent to choosing a single global model that minimizes the average loss. In the other extreme, when $t\to\infty$, we obtain completely local models except for clients that are deemed to have identical distributions, i.e., when $D_{ij}=0$. In the sequel, we show how to choose the similarity parameters in a completely data-driven manner, based on a surrogate for the Wasserstein distance.
3.2 Model dissimilarity and Wasserstein distance
We now introduce the Wasserstein distance, and connect it to the dissimilarities between the ideal models. Suppose the data space $\mathcal{X}$ is equipped with a metric $d_{\mathcal{X}}$. Then the 1-Wasserstein distance (also known as the earth mover distance and Kantorovich–Rubenstein distance) between two distributions $\mu_{i}$ and $\mu_{j}$ is defined as
$W_{1}(\mu_{i},\mu_{j})=\inf_{\pi\in\Pi(\mu_{i},\mu_{j})}{\mathbbmss{E}}_{(\xi,\xi^{\prime})\sim\pi}[d_{\mathcal{X}}(\xi,\xi^{\prime})],$
where $\Pi(\mu_{i},\mu_{j})$ is the set of all probability measures with marginals $\mu_{i}$ and $\mu_{j}$. This distance is indeed a true metric on the space of distributions on $\mathcal{X}$, meaning it is symmetric, positive definite, and satisfies the triangle inequality (Villani,, 2008).
We will now show that under two mild conditions, the dissimilarities of the ideal models for two distributions can be bounded by the 1-Wasserstein distance. The first assumption is that the losses grow at a quadratic rate when diverting from the ideal model. This condition is satisfied by all losses whose convex envelopes are strongly convex. The second assumption is that the loss function is Lipschitz continuous with respect to the data, i.e., a small difference between two samples imply a small difference in the value of the loss function $\ell$ at any given model parameters.
Assumption 3.1.
The loss functions $f_{i}$ have quadratic functional growth with parameter $\gamma>0$, meaning for the ideal model $\theta_{i}^{\star}=\operatornamewithlimits{argmin}_{\theta_{i}}f_{i}(\theta_{i})$, the bound $f_{i}(\theta_{i})-f_{i}(\theta_{i}^{\star})\geq(\gamma/2)\|\theta_{i}-\theta_{i}^{\star}\|^{2}$ holds for every $\theta_{i}\in\Theta$.
Assumption 3.2.
The loss function $\ell$ is $L_{\mathcal{X}}$-Lipschitz continuous in its second argument with respect to the metric $d_{\mathcal{X}}$, meaning $|\ell(\theta,\xi)-\ell(\theta,\xi^{\prime})|\leq L_{\mathcal{X}}d_{\mathcal{X}}(\xi,\xi^{\prime})$ for every $\theta\in\Theta$.
Remark 3.3.
The latter assumption holds naturally for some loss functions, but for others, e.g. linear regression with quadratic loss, we would need to require $\Theta$ to be bounded.
Proposition 3.4.
Suppose Assumptions 3.1 and 3.2 hold. Then the ideal models satisfy
$\|\theta_{i}^{\star}-\theta_{j}^{\star}\|_{2}^{2}\leq\frac{2L_{\mathcal{X}}}{\gamma}W_{1}(\mu_{i},\mu_{j}).$ (2)
Proof.
The proof of this result, and all forthcoming results, is presented in Appendix A. ∎
Why not estimate model dissimilarity directly?
A straight-forward approach to shrink the model dissimilarities would be to learn completely local models first, calculate their pairwise distances, and use them as the dissimilarity parameters. So why use a less direct approach with Wasserstein distances? In the FL setting, it is more common than not to have more features than observations on each client. In, e.g., a linear regression task with $d$ important features and clients with $N$ observations each, it is well known that the the minimax squared estimation error is $O(d/N)$ (Bickel et al.,, 2009; Raskutti et al.,, 2011). Moreover, if $N<d$, then without additional assumptions, no meaningful upper bound on the error exists. With models that have more parameters than there are features, this problem is likely to be further exarcerbated. The Wasserstein distance between two empirical distributions, however, has a rate of $O(N^{-1/d})$ (Wang et al.,, 2021, Example 1). This rate is very slow when increasing $N$, but clearly performs much better than $O(d/N)$ when $N$ is on a similar scale as $d$, or is smaller. (In Appendix TODO, we also show empirically that directly estimating the dissimilarities underperforms our proposed approach.)
Choosing the data metric.
In a supervised learning task, a simple choice of data metric $d_{\mathcal{X}}$ is to concatenate the features and the response, and use, e.g., the Euclidean norm. Another approach is given by Elhussein and Gursoy, (2024), who propose using cosine-similarity on the features and the Hellinger distance on the response. They argue cosine-similarity is more informative in high-dimensional feature spaces due to the concentration of measures phenomenon, and that the Hellinger distance compares densities within their statistical manifold, rather than within Euclidean space.
Calculating the dissimilarity parameters.
With the present motivation in mind, we can choose the dissimilarity parameters to be $W_{1}(\hat{\mu}_{i},\hat{\mu}_{j})$, with $\hat{\mu}_{i}$ being the empirical distribution of the $i$th client. Rakotomamonjy et al., (2023) present an iterative method to approximate the Wasserstein distances in a federated manner, which involes computing an optimal transport plan every iteration. This means that in our setting we would have to compute $n(n-1)/2$ optimal transport plans for a number of iterations, which is expensive both in terms of computation and of communication. We therefore propose to approximate the Wasserstein distances via linear embedding of the data sets, following works of e.g. Kolouri et al., (2020) and Liu et al., (2025).
We introduce a reference data set $\mathcal{D}_{0}=\{\xi_{1}^{0},\dots,\xi_{N_{0}}^{0}\}$ of $N_{0}$ independent samples drawn from a reference distribution $\mu_{0}$. The optimal transport plan between the empirical distributions $\hat{\mu}_{0}$ and $\hat{\mu}_{i}$ is then given by
$\pi_{i}^{\star}=\operatornamewithlimits{argmin}_{\pi\in\Pi(\hat{\mu}_{0},\hat{\mu}_{i})}\sum_{j=1}^{N_{0}}\sum_{k=1}^{N_{i}}\pi_{jk}d_{\mathcal{X}}(\xi_{j}^{0},\xi_{k}^{i}).$
By approximating the Monge map via $M_{i}=N_{0}\pi_{i}^{\star}\mathcal{D}_{i}$, we can define an embedding $\Phi$ by $\Phi(\mathcal{D}_{i})=(M_{i}-\mathcal{D}_{0})/\sqrt{N_{0}}$ that has the property that $\|\Phi(\mathcal{D}_{i})-\Phi(\mathcal{D}_{j})\|_{1}$ defines a metric on the space of data sets, and indeed approximates the Wasserstein distance $W_{1}(\hat{\mu}_{i},\hat{\mu}_{j})$. Thus, we choose $D_{ij}=\|\Phi(\mathcal{D}_{i})-\Phi(\mathcal{D}_{j})\|_{1}$ as the dissimilarity parameters. By doing this, each client only has to compute a single optimal transport plan. Also, on the communcation side, only a reference data set and the embeddings need to be transmitted over the network once, totalling $O(N_{0}dn)$ communication.
3.3 Algorithm
To solve the Karula problem (1), we need an appropriate algorithm. Here, we present a projected gradient algorithm that need only a small subset of clients to participate in each training round. To eliminate the error introduced by partial participation in each training round, our algorithm uses variance reduction, taking inspiration from methods such as Saga (Defazio et al.,, 2014) and Fedvarp (Jhunjhunwala et al.,, 2022). Our algorithm can be viewed as a relaxation of standard distributed stochastic gradient method, where our algorithm in each iteration shrinks the dissimilarities between the local models, instead of setting them to zero, i.e., averaging. Specifically, consider the following formulation of the Karula problem (1):
$\begin{array}[]{ll}\underset{\theta\in\mathcal{K}}{\mbox{minimize}}&f(\theta)=\frac{1}{n}\sum_{i=1}^{n}\alpha_{i}f_{i}(\theta_{i}),\quad\mathcal{K}=\{\theta\in{\mbox{\bf R}}^{d\times n}\colon\|\theta_{i}-\theta_{j}\|^{2}\leq tD_{ij}\}.\end{array}$
Suppose that each client $i$ may function as a stochastic first-order oracle $G_{i}$ that returns a stochastic gradient given an input model. With this formulation we can apply a stochastic projected gradient method, which is presented in Algorithm 1. In each training round of the algorithm, a subset of the clients return a stochastic gradient to the server. Subsequently, the server performs a variance reduced gradient step with respect to $f$, and projects the iterate onto $\mathcal{K}$.
The projection $\Pi_{\mathcal{K}}(\vartheta)$ of a collection of models $\vartheta\in{\mbox{\bf R}}^{d\times n}$ onto $\mathcal{K}$ does not have an analytical solution, so we need to employ an iterative method to compute it. Therefore, we define the $\delta$-inexact projection $\Pi_{\mathcal{K}}^{\delta}\colon\Theta^{n}\to\mathcal{K}$ that returns a feasible collection of models $\theta=\Pi_{\mathcal{K}}^{\delta}(\vartheta)$ which is at most $\delta$-suboptimal in the projection problem, meaning
$\frac{1}{2\eta}\|\theta-\vartheta\|^{2}\leq\delta+\frac{1}{2\eta}\min_{v\in\mathcal{K}}\|v-\theta\|^{2}$
for the stepsize $\eta>0$. We give details on how to solve the projection computation in Appendix TODO.
Algorithm 1 Karula
  Setup: Initializations $(\phi_{1}^{0},\dots,\phi_{n}^{0})$, step size $\eta$, participation size $s$, number of iterations $K$.
  Server receives initializations $\phi_{i}^{0}$ and stochastic gradients $G_{i}(\phi_{i}^{0})$ from clients $i\in\{1,\dots,n\}$, and forms $\theta^{0}=(\phi_{1}^{0},\dots,\phi_{n}^{0})$ and $g^{0}=(G_{1}(\phi_{1}^{0}),\dots,G_{n}(\phi_{n}^{0}))/n$.
  for $k=0,\dots,K-1$ do
     Server samples uniformly a subset $S_{k}\subset\{1,\dots,n\}$ of $s$ clients and sends $\theta_{i}^{k}$ to client $i\in S_{k}$.
     Client $i\in\{1,\dots,n\}$ updates
$\phi_{i}^{k+1}=\begin{cases}\theta_{i}^{k}&\text{if }i\in S_{k},\\ \phi_{i}^{k}&\text{otherwise}.\end{cases}$
     Client $i\in S_{k}$ returns $G_{i}(\phi_{i}^{k+1})$ to the server.
     Server sets
$g_{i}^{k+1}=\begin{cases}G_{i}(\phi_{i}^{k+1})&\text{if }i\in S_{k},\\ g_{i}^{k}&\text{otherwise}.\end{cases}$
and updates $\nu^{k}=g^{k}+(g^{k+1}-g^{k})/s$.
     Server performs inexact projected gradient step $\theta^{k+1}=\Pi_{\mathcal{K}}^{\delta}(\theta^{k}-\eta\nu^{k})$.
  end for
4 Convergence analysis
In this section, we present an analysis of the convergence properties of our proposed algorithm for solving Problem (1), both in the general smooth case when the losses have Lipschitz continuous gradients, and in the smooth and strongly convex case. The analysis is inspired by the work of Reddi et al., (2016), which we have adapted for the FL setting, and to allow for inexact projections and stochastic gradients of each loss $f_{i}$.
In the smooth, possibly nonconvex case, we are interested in finding a local minimizer, but due to the problem being constrained, the ordinary stationarity condition $0=\nabla f(\theta)$ is not suitable. In lieu, the analogous condition for constrained optimization is $\theta$ being a zero of the gradient mapping
$\mathbf{G}_{\eta}(\theta)=\frac{1}{\eta}\left(\theta-\Pi_{\mathcal{K}}(\theta-\eta\nabla f(\theta)\right).$
Informally, we can think of the condition $0=\mathbf{G}_{\eta}(\theta)$ as saying one of two things: either $\theta$ is a feasible point with zero gradient, or it is a point such that the negative gradient is normal to the feasible set.
To establish the convergence results, we make some standard assumptions. First, we assume smoothness of the losses.
Assumption 4.1.
The weighted losses $\alpha_{i}f_{i}$ are $L$-smooth, meaning $\alpha_{i}\|\nabla f_{i}(\theta_{i})-\nabla f_{i}(\theta_{i}^{\prime})\|\leq L\|\theta_{i}-\theta_{i}^{\prime}\|$ for all $\theta_{i}$ and $\theta_{i}^{\prime}$.
Second, we will assume that the stochastic gradients have bounded mean squared error.
Assumption 4.2.
The stochastic gradients returned by the oracles $G_{i}$ have bounded mean squared error, meaning ${\mathbbmss{E}}\|G_{i}(\theta_{i})-\alpha_{i}\nabla f_{i}(\theta_{i})\|^{2}\leq\sigma^{2}$ for every $\theta_{i}$.
Third, we will make use of strong convexity in one of the forthcoming convergence results.
Assumption 4.3.
The losses $f_{i}$ are $\gamma$-strongly convex, meaning $f_{i}(\theta_{i})\leq f_{i}(\theta_{i}^{\prime})+\langle\nabla f_{i}(\theta_{i}),\theta_{i}^{\prime}-\theta_{i}\rangle+(\gamma/2)\|\theta_{i}-\theta_{i}^{\prime}\|^{2}$.
We find under the two former assumptions that the proposed algorithm enjoys sublinear convergence to an $\epsilon$-stationary point, with the error $\epsilon$ arising due to the stochastic gradients and inexact projections.
Theorem 4.4.
Under Assumptions 4.1 and 4.2, with feasible initialization, step size $\eta=3/(8L)$ and $12\leq s\leq n-$ 1, the iterates of Algorithm 1 satisfy
$\min_{k=0,\dots,K-1}{\mathbbmss{E}}\|\mathbf{G}_{\eta}(\theta^{k})\|^{2}\leq\frac{8L}{3}\left(\frac{f(\theta^{0})-f(\theta^{\star})}{K}+\epsilon\right),$
where $\epsilon=4\sigma^{2}(s+1)/s+2\delta$.
The assumption of feasible initialization is only made for the sake of exposition, and is trivially satisfied by initializing all clients to the same model.
Theorem 4.5.
Under Assumptions 4.1, 4.2 and 4.3, with feasible initialization, step size $\eta=3/(8L)$ and $12\leq s\leq n-$ 1,the iterates of Algorithm 1 satisfy for $k_{m}=mR\leq K$,
$\min_{k\in[k_{m},k_{m+1}]}{\mathbbmss{E}}[f(\theta^{k})-f(\theta^{\star})]\leq\frac{f(\theta^{0})-f(\theta^{\star})}{\rho^{m}}+\kappa\epsilon,$
where $\rho=\kappa R$ and $\epsilon=4\sigma^{2}(s+1)/s+2\delta$, and $\kappa=\gamma/L$.
Remark 4.6.
Note that in contrast to many other FL algorithms, our convergence analysis require no bounding of client heterogeneity for the convergence or other assumptions about the distributions, and instead relies only on standard machine learning assumptions. Moreover, the step size that is used to achieve the bound is comparatively genereous, and works well in practice when the smoothness constant $L$ is known or can be effectively approximated.
Remark 4.7.
While it is common to assume that the projections are exact, it is not particularly realistic. With our analysis, we explicitly account for error in the projection and show the contribution of the projection error to the neighborhood that the algorithm converges to. This informs practice, since we can choose the error $\delta$ to trade off the iteration cost with proximity to stationarity.
5 Numerical experiments
Here, we study the performance of Karula empirically and compare to some existing conventional and personalized FL strategies. First, in §5.1, we explore ridge regression on synthetic, heterogeneous data where the true model parameters are known. Then, in §5.2, we compare the different methods on a small-scale logistic regression task on heart disease data distributed accross four hospitals. Last, in §5.3, we evaluate the performance of the different methods on a federated handwritten digit classification task using neural network models.
Benchmark strategies.
We compare Karula with five other FL strategies: local non-collaborative learning, federated stochastic gradient method, federated averaging, pFedMe and IFCA. For all methods, we weight the client loss functions by the number of samples.
5.1 Synthetic data
We begin by comparing the performance of the strategies on a ridge regression problem with $d=50$ features. The ridge penalization parameter is chosen as $\lambda=10^{-6}$, to ensure a unique solution. One third of the clients participate at every round and clients return full gradients.
Data generation.
We generate a network of $n=30$ clients, where the ground-truth model parameters for a third of the clients are drawn from a normal distribution with mean $1$, another third from a normal distribution with mean $1.5$, and the remaining third from a normal distribution with mean $2$. We draw feature matrices with entries from three different normal distributions as well. Each client draws $N_{i}$ samples from their respective distribution, with $N_{i}$ being uniformly distributed between 10 and 100. We then set the underlying noise such that the global model performs slightly better than the completely local model.
Regularization parameter selection.
To set as level of a playing field as possible for the different FL strategies, we rely on 5-fold cross-validation to select the regularization parameters of pFedMe, IFCA and Karula. We do however acknowledge that it is not always feasible to do so in large-scale FL systems in practice.
Results.
Table 1: Summary statistics of the performances of the different strategies on synthetic and heart disease data, with ±$\pm 2\times\mathrm{SE}$.
(a) Mean estimation error and test $R^{2}$ on synthetic data.
Strategy Estimation error Test $R^{2}$
Local 35.33 $\pm$ 2.729 0.692 $\pm$ 0.108
FedSGD 7.60 $\pm$ 0.513 0.839 $\pm$ 0.024
FedAvg 7.47 $\pm$ 0.493 0.846 $\pm$ 0.026
pFedMe 31.65 $\pm$ 0.986 0.768 $\pm$ 0.044
IFCA 7.69 $\pm$ 0.509 0.821 $\pm$ 0.024
Karula 5.86 $\pm$ 0.204 0.938 $\pm$ 0.009
(b) Test accuracy on heart disease data.
Strategy Test accuracy
Local 0.658
FedSGD 0.685
FedAvg 0.669
pFedMe 0.658
IFCA -
Karula 0.705
5.2 Federated heart disease data
We use the heart disease dataset provided by Janosi et al., (1988), which contains data on 940 patients spread out over four different hospitals in the US and in Europe. We preprocess the data as in (Ogier du Terrail et al.,, 2022), giving us 13 features and in total 740 samples, with a range of 46 to 303 patient samples per hospital.
5.3 Federated MNIST
We run experiments on a federated version of the MNIST dataset (Caldas et al.,, 2019). The task is classification, with ten different classes corresponding to digits $0,\dots,9$. Each client in the federated network owns data of handwritten digits corresponding to one individual writer. Here, heterogeneity in the network stems from idiosyncratic handwriting; some write small and others write large, some write “1” with serifs and others write without, etc.
Results.
Figure 1:
6 Discussion
Appendix A Proofs of results
For simplicity, suppose without loss of generality that $\Theta={\mbox{\bf R}}^{p}$ for some natural number $p$.
A.1 Proof of Proposition 3.4
Proof.
Under Assumptions 3.1, the ideal models $\theta_{i}^{\star}$, $i\in\{1,\dots,n\}$, satisfy
$f_{i}(\theta_{j}^{\star})-f_{i}(\theta_{i}^{\star})$ $\geq\frac{\mu}{2}\|\theta_{i}^{\star}-\theta_{j}^{\star}\|_{2}^{2},$
$f_{j}(\theta_{i}^{\star})-f_{j}(\theta_{j}^{\star})$ $\geq\frac{\mu}{2}\|\theta_{i}^{\star}-\theta_{j}^{\star}\|_{2}^{2}.$
Rearranging and adding the two inequalities, we have that
$\|\theta_{i}^{\star}-\theta_{j}^{\star}\|_{2}^{2}\leq\frac{1}{\mu}\left(f_{i}(\theta_{j}^{\star})-f_{j}(\theta_{j}^{\star})+f_{j}(\theta_{i}^{\star})-f_{i}(\theta_{i}^{\star})\right).$ (3)
By Kantorovich–Rubenstein Duality (Villani,, 2008), it holds for all $\theta$ that
$f_{i}(\theta)-f_{j}(\theta)$ $={\mathbbmss{E}}_{\xi\sim\mu_{i}}[\ell(\theta,\xi)]-{\mathbbmss{E}}_{\xi\sim\mu_{j}}[\ell(\theta,\xi)]$
$\leq L_{\mathcal{X}}\sup_{g\in\mathcal{L}_{1}}\left\{{\mathbbmss{E}}_{\xi\sim\mu_{i}}[g(\xi)]-{\mathbbmss{E}}_{\xi\sim\mu_{j}}[g(\xi)]\right\}$
$=L_{\mathcal{X}}W_{1}(\mu_{i},\mu_{j}),$
where Assumption 3.2 was used in the inequality, and where $\mathcal{L}_{1}$ is the space of $1$-Lipschitz functions. Combining this inequality with (3), we obtain the stated bound. ∎
A.2 Technical lemmatta
Lemma A.1.
Under Assumption 4.1 and 4.2, the iterates $\theta^{k}$, $\phi^{k}$ and $\nu^{k}$ of Algorithm 1 satsify
${\mathbbmss{E}}\|\nu^{k}-\nabla f(\theta^{k})\|^{2}\leq\frac{2L^{2}}{ns}{\mathbbmss{E}}\|\theta^{k}-\phi^{k}\|^{2}+\frac{4(s+1)}{s}\sigma^{2}.$
Proof.
First, define $\Delta^{k}=(\Delta_{1}^{k},\dots,\Delta_{n}^{k})$, with
$\Delta_{i}^{k}=\begin{cases}\pi_{i}(\nabla f_{i}(\theta_{i}^{k})-\nabla f_{i}(\phi_{i}^{k}))/s&\text{if }i\in S_{k},\\ 0&\text{otherwise},\end{cases}$
and note that ${\mathbbmss{E}}\Delta^{k}=\nabla f(\theta^{k})-\nabla f(\phi^{k})$. Recall also that $\nu^{k}=g^{k}+(g^{k+1}-g^{k})/s$. Thus, we have
${\mathbbmss{E}}\|\nu^{k}-\nabla f(\theta^{k})\|^{2}$ $\leq{\mathbbmss{E}}\left[2\|\nabla f(\theta^{k})-(\nabla f(\phi^{k})+\Delta^{k}))\|^{2}+2\|\nu^{k}-(\nabla f(\phi^{k})+\Delta^{k})\|^{2}\right]$
$=2{\mathbbmss{E}}\left[\|\Delta^{k}-(\nabla f(\theta^{k})-\nabla f(\phi^{k})))\|^{2}+\|g^{k}+(g^{k+1}-g^{k})/s-(\nabla f(\phi^{k})+\Delta^{k})\|^{2}\right]$
$\leq 2{\mathbbmss{E}}\|\Delta^{k}-{\mathbbmss{E}}\Delta^{k}\|^{2}+\frac{4(s+1)}{s}\sigma^{2}$
$\leq 2{\mathbbmss{E}}\|\Delta^{k}\|^{2}+\frac{4(s+1)}{s}\sigma^{2}$
$=2{\mathbbmss{E}}\left[\frac{1}{s^{2}}\sum_{i\in S_{k}}\alpha_{i}^{2}\left\|\nabla f_{i}(\theta_{i}^{k})-\nabla f_{i}(\phi_{i}^{k})\right\|^{2}\right]+\frac{4(s+1)}{s}\sigma^{2}$
$\leq 2{\mathbbmss{E}}\left[\frac{L^{2}}{s^{2}}\sum_{i\in S_{k}}\|\theta_{i}^{k}-\phi_{i}^{k}\|^{2}\right]+\frac{4(s+1)}{s}\sigma^{2}$
$\leq\frac{2L^{2}}{ns}{\mathbbmss{E}}\|\theta^{k}-\phi^{k}\|^{2}+\frac{4(s+1)}{s}\sigma^{2}.$
In the the second inequality we used the bounded variance of the stochastic gradients (Assumption 4.2) and the Cauchy–Schwarz inequality coupled with Young’s inequality. In the third inequality we used $L$-smoothness of the weighted losses (Assumption 4.1). ∎
Lemma A.2.
Under Assumption 4.1, for $u=\Pi_{\mathcal{K}}^{\delta}(\theta-\eta\nu)$ and $\bar{u}=\Pi_{\mathcal{K}}(\theta-\eta\nu)$, with any $\alpha>0$, and $\theta,\nu\in{\mbox{\bf R}}^{p}$, the inequality
$f(u)$ $\leq f(v)+\langle u-v,\nabla f(\theta)-\nu\rangle+\left(\frac{L}{2}-\frac{1}{2\eta}\right)\|u-\theta\|^{2}$ (4)
$+\left(\frac{L}{2}+\frac{1}{2\eta}\right)\|v-\theta\|^{2}-\frac{1}{2\eta}\|\bar{u}-v\|^{2}+\delta$
holds for all $v\in\mathcal{K}$. Additionally,
$\frac{1}{2\eta}\|u-\bar{u}\|^{2}\leq\delta.$ (5)
Proof.
By $L$-smoothness of $f$, the two inequalities
$f(u)\leq f(\theta)+\langle\nabla f(\theta),u-\theta\rangle+\frac{L}{2}\|u-\theta\|^{2}$
and
$f(\theta)\leq f(v)+\langle\nabla f(\theta),\theta-v\rangle+\frac{L}{2}\|v-\theta\|^{2}$
hold for any $v\in{\mbox{\bf R}}^{p}$. Adding the two inequalities, we have that
$f(u)\leq f(v)+\langle\nabla f(\theta),u-v\rangle+\frac{L}{2}(\|u-\theta\|^{2}+\|v-\theta\|^{2}).$ (6)
Moreover, since $1/(2\eta)\|v-(\theta-\eta\nu)\|^{2}$ is $1/(2\eta)$-strongly convex in $v$, for all $v\in\mathcal{K}$ the following inequality holds:
$\frac{1}{2\eta}\|\bar{u}-(\theta-\eta\nu)\|^{2}+\frac{1}{2\eta}\|\bar{u}-v\|^{2}\leq\frac{1}{2\eta}\|v-(\theta-\eta\nu)\|^{2}.$
We also have that
$\frac{1}{2\eta}\|u-(\theta-\eta\nu)\|^{2}-\delta\leq\frac{1}{2\eta}\|\bar{u}-(\theta-\eta\nu)\|^{2},$ (7)
so,
$\frac{1}{2\eta}\|u-(\theta-\eta\nu)\|^{2}+\frac{1}{2\eta}\|\bar{u}-v\|^{2}\leq\frac{1}{2\eta}\|v-(\theta-\eta\nu)\|^{2}+\delta.$
After some rearranging, we obtain the inequality
$0\leq-\langle u-v,\nu\rangle+\frac{1}{2\eta}(\|u-\theta\|^{2}-\|v-\theta\|^{2}-\|\bar{u}-v\|^{2})+\delta.$
Adding this inequality to inequality (6), we arrive at the first result (4).
For the second result (5), note that due to strong convexity,
$\frac{1}{2\eta}\|\bar{u}-(\theta-\eta\nu)\|^{2}+\frac{1}{2\eta}\|\bar{u}-u\|^{2}\leq\frac{1}{2\eta}\|u-(\theta-\eta\nu)\|^{2}.$
Using this coupled with (7), we obtain the stated inequality.
A.3 Proofs of Theorems
Proof of Theorem 4.4.
Let $\vartheta^{k+1}=\Pi_{\mathcal{K}}(\theta^{k}-\eta\nabla f(\theta^{k}))$ be the full projected gradient update and $\bar{\theta}^{k+1}=\Pi_{\mathcal{K}}(\theta^{k}-\eta\nu^{k})$ be the exact projected update. Using (4) of Lemma A.2, we have that with $u=\theta^{k+1}$, $v=\vartheta^{k+1}$ and $\nu=\nu^{k}$,
${\mathbbmss{E}}[f(\theta^{k+1})]\leq{\mathbbmss{E}}\bigg{[}$ $f(\vartheta^{k+1})+\langle\theta^{k+1}-\vartheta^{k+1},\nabla f(\theta^{k})-\nu^{k}\rangle+\left(\frac{L}{2}-\frac{1}{2\eta}\right)\|\theta^{k+1}-\theta^{k}\|^{2}$ (8)
$+\left(\frac{L}{2}+\frac{1}{2\eta}\right)\|\vartheta^{k+1}-\theta^{k}\|^{2}-\frac{1}{2\eta}\|\bar{\theta}^{k+1}-\vartheta^{k+1}\|^{2}+\delta\bigg{]},$
and with $u=\vartheta^{k+1}$, $v=\theta^{k}$ and $\nu=\nabla f(\theta^{k})$,
${\mathbbmss{E}}[f(\vartheta^{k+1})]\leq{\mathbbmss{E}}\left[f(\theta^{k})+\left(\frac{L}{2}-\frac{1}{2\eta}\right)\|\vartheta^{k+1}-\theta^{k}\|^{2}-\frac{1}{2\eta}\|\vartheta^{k+1}-\theta^{k}\|^{2}\right].$ (9)
To bound the inner product, we apply the Cauchy-Schwarz inequality and Young’s inequality as follows:
${\mathbbmss{E}}[\langle\theta^{k+1}-\vartheta^{k+1},\nabla f(\theta^{k})-\nu^{k}\rangle]$ $\leq{\mathbbmss{E}}\left[\|\theta^{k+1}-\vartheta^{k+1}\|\|\nabla f(\theta^{k})-\nu^{k}\|\right]$
$\leq{\mathbbmss{E}}\left[\frac{1}{4\eta}\|\theta^{k+1}-\vartheta^{k+1}\|^{2}+\eta\|\nabla f(\theta^{k})-\nu^{k}\|^{2}\right]$
$\leq{\mathbbmss{E}}\left[\frac{1}{2\eta}(\|\bar{\theta}^{k+1}-\vartheta^{k+1}\|^{2}+\|\bar{\theta}^{k+1}-\theta^{k+1}\|^{2})+\eta\|\nabla f(\theta^{k})-\nu^{k}\|^{2}\right]$
$\leq{\mathbbmss{E}}\left[\frac{1}{2\eta}\|\bar{\theta}^{k+1}-\vartheta^{k+1}\|^{2}+\eta\|\nabla f(\theta^{k})-\nu^{k}\|^{2}+\delta\right],$
where in the last inequality, we applied Equation (5) from Lemma A.2. Applying Lemma A.1, we obtain the bound
${\mathbbmss{E}}[\langle\theta^{k+1}-\vartheta^{k+1},\nabla f(\theta^{k})-\nu^{k}\rangle]\leq\frac{1}{2\eta}{\mathbbmss{E}}\|\bar{\theta}^{k+1}-\vartheta^{k+1}\|^{2}+\frac{2\eta L^{2}}{ns}{\mathbbmss{E}}\|\theta^{k}-\phi^{k}\|^{2}+\frac{4(s+1)}{s}\sigma^{2}+\delta.$
Adding the inequalities (8) and (9) and using the bound we just derived, we have that
${\mathbbmss{E}}[f(\theta^{k+1})]\leq{\mathbbmss{E}}\bigg{[}$ $f(\theta^{k})+\frac{1}{2\eta}\|\theta^{k+1}-\vartheta^{k+1}\|^{2}+\frac{2\eta L^{2}}{ns}\|\theta^{k}-\phi^{k}\|^{2}+\frac{4(s+1)}{s}\sigma^{2}+2\delta$ (10)
$+\left(L-\frac{1}{2\eta}\right)\|\vartheta^{k+1}-\theta^{k}\|^{2}+\left(\frac{L}{2}-\frac{1}{2\eta}\right)\|\theta^{k+1}-\theta^{k}\|^{2}-\frac{1}{2\eta}\|\theta^{k+1}-\vartheta^{k+1}\|^{2}\bigg{]}$
$={\mathbbmss{E}}\bigg{[}$ $f(\theta^{k})+\frac{2\eta L^{2}}{ns}\|\theta^{k}-\phi^{k}\|^{2}+\frac{4(s+1)}{s}\sigma^{2}+2\delta$
$+\eta^{2}\left(L-\frac{1}{2\eta}\right)\|\mathbf{G}_{\eta}(\theta^{k})\|^{2}+\left(\frac{L}{2}-\frac{1}{2\eta}\right)\|\theta^{k+1}-\theta^{k}\|^{2}\bigg{]}.$
where we used that $\mathbf{G}_{\eta}(\theta^{k})=(\theta^{k}-\vartheta^{k+1})/\alpha$. Now, we introduce the Lyapunov functions
$\nu^{k}={\mathbbmss{E}}\left[f(\theta^{k})+c_{k}\|\theta^{k}-\phi^{k}\|^{2}\right].$
where $c_{k}$ will be set in the sequel. Note that
${\mathbbmss{E}}\|\theta^{k}-\phi^{k}\|^{2}=\frac{s}{n}{\mathbbmss{E}}\|\theta^{k}-\theta^{k-1}\|^{2}+\frac{n-s}{n}{\mathbbmss{E}}\|\theta^{k}-\phi^{k-1}\|^{2},$
hence
$V^{k+1}={\mathbbmss{E}}\bigg{[}$ $f(\theta^{k+1})+c_{k+1}\frac{s}{n}\|\theta^{k+1}-\theta^{k}\|^{2}+c_{k+1}\frac{n-s}{n}\|\theta^{k+1}-\phi^{k}\|^{2}\bigg{]}$
$={\mathbbmss{E}}\bigg{[}$ $f(\theta^{k+1})+c_{k+1}\frac{s}{n}\|\theta^{k+1}-\theta^{k}\|^{2}+c_{k+1}\frac{n-s}{n}\|(\theta^{k+1}-\theta^{k})-(\phi^{k}-\theta^{k})\|^{2}\bigg{]}$
$={\mathbbmss{E}}\bigg{[}$ $f(\theta^{k+1})+c_{k+1}\frac{s}{n}\|\theta^{k+1}-\theta^{k}\|^{2}$
$+c_{k+1}\frac{n-s}{n}(\|\theta^{k+1}-\theta^{k}\|^{2}+2\langle\theta^{k+1}-\theta^{k},\theta^{k}-\phi^{k}\rangle+\|\theta^{k}-\phi^{k}\|^{2})\bigg{]}$
$={\mathbbmss{E}}\bigg{[}$ $f(\theta^{k+1})+c_{k+1}\|\theta^{k+1}-\theta^{k}\|^{2}+c_{k+1}\frac{n-s}{n}(2\langle\theta^{k+1}-\theta^{k},\theta^{k}-\phi^{k}\rangle+\|\theta^{k}-\phi^{k}\|^{2})\bigg{]}.$
Once again, we use the Cauchy–Schwarz inequality and Young’s inequality to bound the inner product:
${\mathbbmss{E}}\left[\langle\theta^{k+1}-\theta^{k},\theta^{k}-\phi^{k}\rangle\right]\leq{\mathbbmss{E}}\left[\frac{n}{2s}\|\theta^{k+1}-\theta^{k}\|^{2}+\frac{s}{2n}\|\theta^{k}-\phi^{k}\|^{2}\right].$
Combining this with inequality (10), it follows that
$V^{k+1}\leq{\mathbbmss{E}}\bigg{[}$ $f(\theta^{k+1})+c_{k+1}\frac{n}{s}\|\theta^{k+1}-\theta^{k}\|^{2}+c_{k+1}\frac{n-s}{n}\left(1+\frac{s}{n}\right)\|\theta^{k}-\phi^{k}\|^{2}\bigg{]}$
$\leq{\mathbbmss{E}}\bigg{[}$ $f(\theta^{k})+\frac{2\eta L^{2}}{ns}\|\theta^{k}-\phi^{k}\|^{2}+\frac{4(s+1)}{s}\sigma^{2}+2\delta$
$+\eta^{2}\left(L-\frac{1}{2\eta}\right)\|\mathbf{G}_{\eta}(\theta^{k})\|^{2}+\left(\frac{L}{2}-\frac{1}{2\eta}\right)\|\theta^{k+1}-\theta^{k}\|^{2}$
$+c_{k+1}\frac{n}{s}\|\theta^{k+1}-\theta^{k}\|^{2}+c_{k+1}\frac{n-s}{n}\left(1+\frac{s}{n}\right)\|\theta^{k}-\phi^{k}\|^{2}\bigg{]}$
$={\mathbbmss{E}}\bigg{[}$ $f(\theta^{k})+\eta^{2}\left(L-\frac{1}{2\eta}\right)\|\mathbf{G}_{\eta}(\theta^{k})\|^{2}+\left(\frac{L}{2}-\frac{1}{2\eta}+c_{k+1}\frac{n}{s}\right)\|\theta^{k+1}-\theta^{k}\|^{2}$
$+\left(\frac{2\eta L^{2}}{ns}+c_{k+1}\frac{n^{2}-s^{2}}{n^{2}}\right)\|\theta^{k}-\phi^{k}\|^{2}+\frac{4(s+1)}{s}\sigma^{2}+2\delta\bigg{]}.$
Now, setting $c_{K}=0$ and recursively defining
$c_{k}=\frac{2\eta L^{2}}{ns}+c_{k+1}\frac{n^{2}-s^{2}}{n^{2}}$
it holds that
$c_{k}\leq c_{0}=\frac{2\eta L^{2}}{ns}\frac{1-\left(\frac{n^{2}-s^{2}}{n^{2}}\right)^{K}}{\frac{n^{2}-s^{2}}{n^{2}}}\leq\frac{2\eta L^{2}n}{s(n^{2}-s^{2})}$
for all $k=1,\dots,K$, hence
$\frac{L}{2}-\frac{1}{2\eta}+c_{k+1}\frac{n}{s}\leq\frac{L}{2}-\frac{1}{2\eta}+\frac{2\eta L^{2}n^{2}}{s^{2}(n^{2}-s^{2})}.$
When is the right-hand side nonpositive? Consider the quantity $q(s^{2})=s^{2}(n^{2}-s^{2})/n^{2}$, and note that it is concave in $s^{2}\in[2^{2},(n-1)^{2}]$, meaning its minimum is attained at one of the end points. But for $s\in\{2,n-1\}$, $q(s^{2})>1$ holds for all $n\geq 3$. Hence,
$\frac{L}{2}-\frac{1}{2\eta}+\frac{2\eta L^{2}n^{2}}{s^{2}(n^{2}-s^{2})}\leq\frac{L}{2}-\frac{1}{2\eta}+2\eta L^{2}<0$
for $\eta=3/(8L)$. Thus,
$V^{k+1}$ $\leq{\mathbbmss{E}}\left[f(\theta^{k})+\eta^{2}\left(L-\frac{1}{2\eta}\right)\|\mathbf{G}_{\eta}(\theta^{k})\|^{2}+c_{k}\|\theta^{k}-\phi^{k}\|^{2}+\frac{4(s+1)}{s}\sigma^{2}+2\delta\right]$
$=\nu^{k}+\left(L-\frac{1}{2\eta}\right){\mathbbmss{E}}\|\mathbf{G}_{\eta}(\theta^{k})\|^{2}+\frac{4(s+1)}{s}\sigma^{2}+2\delta.$
Summing over $k\in\{0,\dots,K-1\}$, and using the fact that $\nu^{k}=f(\theta^{K})$ due to $c_{K}=0$ and $V^{0}=f(\theta^{0})$ due to $\theta^{0}=\phi^{0}$, we arrive at the inequality
$f(\theta^{K})\leq f(\theta^{0})+K\left(\frac{4(s+1)}{s}\sigma^{2}+2\delta\right)+\sum_{k=0}^{K-1}\eta\left(L-\frac{1}{2\eta}\right){\mathbbmss{E}}\|\mathbf{G}_{\eta}(\theta^{k})\|^{2}.$
Rearranging and dividing by $K$,
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbbmss{E}}\|\mathbf{G}_{\eta}(\theta^{k})\|^{2}$ $\leq\frac{1}{\frac{\eta}{2}-\eta^{2}L}\left(\frac{f(\theta^{0})-f(\theta^{K})}{K}+\frac{4(s+1)}{s}\sigma^{2}+2\delta\right)$
$\leq\frac{2}{\eta(1-2\eta L)}\left(\frac{f(\theta^{0})-f(\theta^{\star})}{K}+\frac{4(s+1)}{s}\sigma^{2}+2\delta\right).$
Lastly, lower bounding the terms in the left hand side by the smallest $\|\mathbf{G}_{\eta}(\theta^{k})\|^{2}$, we arrive at the stated result. ∎
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [TODO]
[…]
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [TODO]
[…]
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [TODO]
[…]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [TODO]
[…]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [TODO]
[…]
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [TODO]
[…]
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [TODO]
[…]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [TODO]
[…]
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [TODO]
[…]
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [TODO]
[…]
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [TODO]
[…]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [TODO]
[…]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [TODO]
[…]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [TODO]
[…]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [TODO]
[…]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [TODO]
[…]
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.